How much imbalance of protons to electrons numbers have to be done to the bulk Earth and Moon in order to make the electrostatic interaction dominate the gravity by a bare minimum magnitude?
1
Expert's answer
2020-09-02T13:24:25-0400

There should be an imbalance of "3.55\u00d710^{32}" number of charges on moon as well as on Earth

. Explanation

"G=6.67\u00d710^{-11}Nm^{2}Kg^{24}"

"Mass" "Of" "Earth(M)=5.97\u00d710^{24}Kg"

"Mass" "Of" "Moon(m)=7.3\u00d710^{22}Kg"

mean distance between earth and moon "d=3.84\u00d710^8m"


Gravitational force between earth and moon is given as.

"\implies F_G=\frac{GMm}{d^{2}}"

"\implies=\frac{(6.67×10^{-11})×(5.97×10^{24})×(7.3×10^{22})}{(3.84×10^{8})^2}"


"F_G=1.97\u00d710^{20}N"


Let there be n+be charges on Earth and n+be charges on moon

so charge of earth "Q_e=n\u00d71.6\u00d710^{-19}"

charge on moon "Q_m=n\u00d71.6\u00d710^{-19}C"

"K=9\u00d710^{9}Nm^{2}C^{-2}"


Then the electrostatic force between the earth and moon

. "F_E=\frac{KQ_mQ_e}{d^{2}}"

"F_E=\frac{(9.9×10^{9})×(n×1.6×10^{-19})×(n×1.6×10^{-19}}{(3.84×10^{8})^2}."


For electrostatic interaction to be just greater than the gravitational interaction,the magnitude of the gravitational force and electrostatic force should be equal

"\implies \mid F_E\mid =\mid F_G\mid"

"\implies n^{2}×1.56×10^{-45}=1.97×10^{20}N"

"\implies n^{2}=1.26×10^{65}"

"n=3.55\u00d710^{32}"

So there should be an imbalance of "3.55\u00d710^{32}" Number of charges on moon as well as on Earth